OpenStudy (anonymous):

Factor the expression, x^2+6x+5

Directrix (directrix):

What are two numbers that multiply to 5 AND add to 6?

OpenStudy (anonymous):

Hmm, 1 and 5?

Directrix (directrix):

x^2+6x+5 = x^2 + 5x + 1x + 5 = x*(x + 5) + 1*(x + 5)= (x + 5)*(x + 1)

Directrix (directrix):

Notice that the 6x was split into 1x + 5x for factoring by grouping. Then, there was a common factor of (x + 5) to factor out, leaving (x + 1)

OpenStudy (anonymous):

Okay lemme write this on paper one sec :)

OpenStudy (anonymous):

how did you go from x*(x + 5) + 1*(x + 5)= to (x + 5)*(x + 1)

OpenStudy (anonymous):

You're multiplying both x and 1 by (x+5), so you can rewrite it. Distributive property I think.

Directrix (directrix):

Consider: 5x + 5 y = 5*(x + y) where 5 is a common monomial factor. Think about the (x + 5) in the same way and factor it out.